# A Spatio-Temporal Machine Learning Model for Mortgage Credit Risk: Default Probabilities and Loan Portfolios

### v1

@@ 3.1 Data and default definition
In total, our data set contains 2’256’528 loan-year observations for 538’942 different mortgages, of which 35’923 loans defaulted. Figure 1 shows the number of defaults and the default rate over time. Many defaults occur around the year 2009 during the global financial crisis, and a particular one-year spike occurs in the year 2020 during the COVID-19 pandemic.
@@
For the single-family loan-level data set, Freddie Mac provides the first three digits of a five-digit postal code for the property of each mortgage. For confidentiality reasons, exact locations are not available. Every three-digit postal code is associated with a specific area in the United States, and we assign the centroid coordinates of the corresponding area to every mortgage. Our data set contains mortgages from a total of 875 different areas. Figure 2 shows the aggregate default rates for the three-digit postal code areas over the years 2000 to 2022. Spatial patterns are visible such as higher default rates in the states of California, Florida, Nevada, and Arizona, and lower default rates in states such as Idaho, Wyoming, North Dakota, and Nebraska.
@@ 3.2 Predictor variables
From the data provided by Freddie Mac, we construct several static and temporally varying predictor variables. The latter predictor variables are based on information available at the beginning of each year and include the age of the loan in months, the current loan-to-value ratio, and the interest rate spread. The current loan-to-value ratio is calculated by dividing the current unpaid principal balance by the appraised value of the property at the time of origination. To calculate the interest rate spread, we follow Calabrese et al. (2024) and subtract from the loan interest rates the average interest rate currently issued by Freddie Mac for 30-year fixed-rate mortgages. This average interest rate is retrieved from the Federal Reserve Bank of St. Louishttps://fred.stlouisfed.org/series/MORTGAGE30US. In addition, we use year-fixed effects. Static predictor variables include the borrower’s credit score at the time when the mortgage originated, additional financial indicators, and several characteristics about the property and the mortgage. A description of all predictor variables is provided in Table 1. These predictor variables are commonly chosen when Freddie Mac’s loan-level data set is used in the credit risk literature; see, e.g., Hu and Zhou (2019) and Medina-Olivares et al. (2023a). Missing values for categorical predictor variables are imputed using the most frequent category, and numeric predictor variables are imputed using the mean. Summary statistics of the numeric and categorical variables are provided in Tables 5 and 6, respectively, in Appendix A.5. For all linear models, we additionally include an intercept term in the predictor variables $X_{ki}$.
@@
orig_dti Original dept-to-income: borrower’s monthly debt payments divided by the borrower’s monthly income
orig_cltv Original combined loan-to-value: original mortgage loan plus any secondary mortgage loan divided by the mortgage appraised value
orig_upb Original unpaid principal balance of the mortgage
@@
year_versioning Year of the loan-year observation
cnt_ltv Current loan-to-value: current unpaid principal balance divided by the mortgage appraised value
@@ 3.3 Models considered and implementation details
We consider the following models with increasing levels of complexity: an independent linear hazard model, a linear spatial model, a linear spatio-temporal model, and a tree-boosted spatio-temporal frailty model; see Section 2. For the spatial and spatio-temporal Gaussian process models, we use a Matérn covariance function as defined in Equations (4) and (5) with smoothness parameter $\nu=1.5$ and $m=20$ nearest neighbors for estimation and prediction with Vecchia approximations. For purely spatial processes, we use the Euclidean distance to determine the nearest neighbors in Vecchia approximations and a random ordering of the spatial coordinates since this gives accurate approximations (Guinness, 2018). For the spatio-temporal models, we use a correlation-based approach to determine nearest neighbors as in Kang and Katzfuss (2023) and an increasing temporal order and a random spatial order for coordinates with the same time. Following Kang and Katzfuss (2023), the nearest neighbors for the spatio-temporal models are redetermined in every iteration that is a power of two whenever an optimization algorithm is used for determining the parameters $\theta$. For prediction with Vecchia approximations, the observed points appear first in the ordering, and the Gaussian process at prediction points is only conditioned on the training data in the Vecchia approximation. For finding optima for the parameters $\theta$, we use the limited-memory BFGS algorithm for the linear Gaussian process models and gradient descent with Nesterov acceleration for the tree-boosted spatio-temporal frailty model. Estimation and prediction with the linear spatial, linear spatio-temporal, and the tree-boosted spatio-temporal frailty model is done using the GPBoosthttps://github.com/fabsig/GPBoost library version 1.4.0 (Sigrist et al., 2021). The code for executing the mortgage credit risk application and generating the data set is publicly available; see https://github.com/pkuendig/SpaceTimeFrailty.
@@ 3.4 Sample split for model evaluation and choosing tuning parameters
We conduct one-year-ahead default predictions for each year starting from 2008 through 2022 using an expanding window training data approach. I.e., we first consider all loan-year observations up to and including the year 2007 as training data and make predictions for all loan-year observations of 2008. We then continue by expanding the training window by one year and using the subsequent year as test data. I.e., all the following results are based on temporal out-of-sample predictions.
For choosing tuning parameters, we analogously split every training data set into an inner training data set and a validation data set. The validation data contains all loan-year observations of the most recent year in the training data, and the inner training data consists of all samples excluding this most recent year. Tuning parameters are chosen by estimating models on the inner training data and selecting the combinations of tuning parameters that maximize the area under the receiver operating characteristic curve (AUC) on the validation data. The candidate tuning parameters for the tree-boosted spatio-temporal frailty model are shown in Table 4 in Appendix A.4.
@@ 3.5 Prediction of individual default probabilities
Table 2 reports the AUC, the H-measure, mean log-loss, Brier score, and ECE for every model, averaged over the 15 years for which we perform one-year-ahead default predictions. We find that the predictive default probabilities of the tree-boosted spatio-temporal frailty model are the most accurate for all metrics considered. An independent linear hazard model has overall the worst prediction accuracy. Adding a spatial frailty variable to an independent linear hazard model improves the prediction accuracy. Further, the linear spatio-temporal model outperforms a purely spatial linear model in all prediction accuracy measures. We interpret this in the sense that there are both non-linearities and/or interactions among the predictor variables as well as spatio-temporal frailty effects present in the mortgage data. We corroborate this conclusion when analyzing the estimated models in more detail in Section 3.7.
@@
Linear independent 0.7649 0.1951 $8.101\times 10^{-2}$ $1.682\times 10^{-2}$ $1.177\times 10^{-2}$
Linear spatial 0.7664 0.1981 $8.092\times 10^{-2}$ $1.686\times 10^{-2}$ $1.188\times 10^{-2}$
Linear spatio-temporal 0.7711 0.2079 $8.032\times 10^{-2}$ $1.679\times 10^{-2}$ $1.164\times 10^{-2}$
LaGaBoost spatio-temporal 0.7769 0.2148 $\mathbf{7.859\times 10^{-2}}$ $\mathbf{1.638\times 10^{-2}}$ $\mathbf{1.017\times 10^{-2}}$
@@
Figure 3 additionally reports the AUC of every model over time. We observe that the tree-boosted spatio-temporal frailty model has the highest AUC for most years. Further, the AUC is at a relatively high level for all models and for all years before the COVID-19 pandemic. In particular, the AUC remains at constant high levels during the global financial crisis around the year 2009. However, in the year 2020 during the COVID-19 pandemic, the AUC drops to considerably lower levels for all models and increases again in the subsequent years. Similar patterns over time are observed for the H-measure and the ECE; see Figures 10 and 11 in Appendix A.1. We interpret this sudden decrease in prediction accuracy of all models in 2020 in the sense that an artificial, external shock in the form of lock-downs led to very different default mechanisms compared to normal times and also compared to the global financial crisis around the year 2009.
Figure 3: Temporal out-of-sample test area under the receiver operator curve (AUC) (higher = better).
@@ 3.6 Prediction of loan portfolio loss distributions
Table 3 reports the CRPS, the 99% quantile loss, and the RMSE for the 15 years used as test data. We observe that the tree-boosted spatio-temporal model has the lowest CRPS, quantile loss, and RMSE compared to all other models. I.e., predictive portfolio loss distributions of the tree-boosted spatio-temporal model are the most accurate when looking at the entire distribution, upper quantiles, and the center of the distribution. Further, the predictive portfolio loss distributions of the linear spatio-temporal model are more accurate in terms of the CRPS and the 99% quantile loss compared to the linear spatial model and also the independent linear hazard model. The linear spatial model has a lower 99% quantile loss but a higher RMSE and CRPS than the independent linear model. I.e., as expected, adding a spatial or a spatio-temporal frailty variable improves the accuracy of upper tail predictions compared to a linear model assuming independence conditional on observable predictor variables.
@@
Linear independent $2.873\times 10^{8}$ $1.237\times 10^{8}$ $5.082\times 10^{8}$
Linear spatial $2.882\times 10^{8}$ $1.229\times 10^{8}$ $5.088\times 10^{8}$
Linear spatio-temporal $2.790\times 10^{8}$ $1.125\times 10^{8}$ $5.148\times 10^{8}$
LaGaBoost spatio-temporal $\mathbf{2.553\times 10^{8}}$ $\mathbf{1.058\times 10^{8}}$ $\mathbf{4.710\times 10^{8}}$
@@
Next, Figure 4 shows the differences between the means of the predictive portfolio loss distributions and the realized portfolio losses over the years 2008 to 2014. Interestingly, among the models considered, the tree-boosted spatio-temporal frailty model predicts the highest and most realistic mean portfolio loss at the beginning of the global financial crisis during the years 2008 and 2009, and subsequently, its mean portfolio loss decreases faster to more realistic levels after the crisis. From a risk management perspective, such a lower pre-crises underestimation and lower post-crisis overestimation of portfolio loss is clearly desirable. In Figure 12 in Appendix A.2, we also show the differences between the predictive mean portfolio losses and the realized portfolio losses for all prediction years from 2008 to 2022.
@@
Figure 5 additionally shows the time series of predictive $99\%$ quantiles of one-year-ahead portfolio loss distributions. Similar to the mean portfolio loss, we find that, among the models considered, the tree-boosted spatio-temporal model predicts the highest upper tail loan portfolio loss at the beginning of the global financial crisis and its predictive upper tail losses are the smallest after the crisis. The independent model without a frailty Gaussian process generates the smallest predictive $99\%$ quantiles before the global financial crisis. The sudden one-year default spike in 2020 during the COVID-19 pandemic causes all classifiers to predict high upper tail loan portfolio losses for the following year 2021.
@@ 3.7 Model interpretation
In the following, we aim to better understand the functioning of the tree-boosted and linear spatio-temporal frailty models. In Figure 6, we show the mean of the posterior distribution of the latent spatio-temporal frailty Gaussian process for the tree-boosted model when training on data up to and including the year 2021, which corresponds to the model for predicting defaults in the most recent year 2022. Due to space constraints, posterior means are not shown for the year 2000. We additionally show these posterior mean maps for the linear spatio-temporal model in Figure 13 in Appendix A.3. We observe that there is considerable variation in the latent frailty variable over space and this variation changes over time. In particular, for the years 2005 and 2006, posterior means for the latent Gaussian variable are high in the New Orleans area, which can be explained by high default rates following Hurricane Katrina in 2005. In 2008 and subsequent years, posterior means are high in the so-called “bubble” states California, Florida, Nevada, and Arizona, where house prices rose particularly rapidly in the run-up to the subprime mortgage crisis (Haughwout et al., 2011). Comparing posterior means of the tree-boosted spatio-temporal frailty model with those of the linear spatio-temporal model, we find similar patterns over space and time.
@@
In Figure 7, we report the estimated covariance parameters for the different expanding window training data sets for the linear spatial, the linear spatio-temporal, and the tree-boosted spatio-temporal frailty models. For the latter, estimated variance and range parameters are relatively constant until the year 2020, when the variance parameter increases from 0.46 to 1.33, and in the following year 2021, the range parameter for time decreases from 2.94 to 1.05. These estimates are likely a consequence of the sudden atypical one-year default spike in 2020 during the COVID-19 pandemic. Since default mechanisms were likely very different in this year compared to the past due to an artificial external shock, less variation in the response variable is explained by the fixed effects and the latent Gaussian variable becomes more important as can be inferred from the higher marginal variance. Subsequently, defaults return to normal levels in 2021, and the estimated correlation with previous years is lower when including 2020 in the training data as can be seen from the lower estimated range parameter for time.
@@
For understanding the function $F(\cdot)$ of the tree-boosted spatio-temporal frailty model, we use SHAP values and SHAP dependence plots (Lundberg and Lee, 2017). Specifically, we consider SHAP values for the model trained on data up to the year 2013. We chose this training window because for the corresponding prediction year 2014, there are pronounced differences in the accuracy of predictive default probabilities between the tree-boosted and linear spatio-temporal frailty models. SHAP values are calculated using 10’000 randomly selected instances of the training data. This subsampling is done to reduce the computational complexity and to avoid that SHAP dependence plots are overcrowded. Using a larger random subsample or a subsample chosen with a different random number generator seed gives almost identical results (results not reported). Figure 8 shows the SHAP values. The predictor variables are ordered according to the average of the absolute values of the SHAP values. According to these results, the eight most important predictor variables in descending order are the credit score (credit_score), the interest rate spread (ir_spread), the current loan-to-value (cnt_ltv), the indicator whether multiple borrowers are obligated to repay (multiple_borrowers), the age of the loan (n_months), the indicator whether the mortgage is used for purchase (loan_purpose = P), the original debt-to-income (orig_dti), and the original unpaid principal balance (orig_upb). For the six numeric variables thereof, we report the corresponding SHAP dependence plots in Figure 9. In all dependence plots, we observe strong interaction effects. This can be inferred from the large vertical scatter of the SHAP values and the systematic colored relationship with other predictor variables. For example, the slope of the SHAP values for the credit score variable (credit_score) is less negative if the interest rate spread variable (ir_spread) is positive. I.e., when interest rate spreads are high, the individual credit score matters less compared to when interest rate spreads are low. In general, the interaction effects are particularly pronounced for very large and small values at the boundaries of the predictor variables. Further, we find strong non-linear effects for the age of the loan (n_months) and the original unpaid principal balance (orig_upb). For instance, the age of a loan is positively related to the default probability up to a certain age of approximately three years, and then the effect flattens out and even starts to slightly decrease. In addition, the variables interest rate spread (ir_spread), current loan-to-value (cnt_ltv), and original debt-to-income (orig_dti) also have clearly non-linear relationships with slopes that change markedly for large and small values of these variables. For instance, increasing interest rate spreads are related to higher default probabilities, but this effect only holds up to a certain level of approximately three percent after which the relationship levels off and further increases in interest rate spreads are not associated with higher default probabilities. Furthermore, the debt-to-income (orig_dti) appears to have an approximately logistic-shaped effect on the default probability with the relationship being almost flat for both small debt-to-income values below 20 percent and large values above 50 percent, and in between, the effect of debt-to-income is approximately linear with a slope that interacts, among other things, with the loan age. We conclude that the presence of interaction and non-linear effects is likely the reason why the tree-boosted spatio-temporal frailty model outperforms the linear spatio-temporal model in terms of prediction accuracy.
@@
Figure 8: SHAP values for the tree-boosted spatio-temporal frailty model when training on data up to the year 2013.
@@
Figure 9: SHAP dependence plots for the tree-boosted spatio-temporal frailty model when training on data up to the year 2013.

### v2

@@ 3.1 Data and default definition
In total, our data set contains 2’256’528 loan-year observations for 538’942 different mortgages, of which 35’923 loans defaulted. Figure 1 shows the number of defaults and the default rate over time. Many defaults occur around the year 2008 during the global financial crisis, and a peculiar one-year spike occurs in the year 2020 during the COVID-19 pandemic.
@@
For the single-family loan-level data set, Freddie Mac provides the first three digits of a five-digit postal code for the property of each mortgage. For confidentiality reasons, exact locations are not available. Every three-digit postal code is associated with a specific area in the United States, and we assign the centroid coordinates of the corresponding area to every mortgage. Our data set contains mortgages from a total of 875 different areas. Figure 2 shows the aggregate default rates for the three-digit postal code areas over the years 2000 to 2022. Spatial patterns are visible such as higher default rates in the states of California, Florida, Nevada, and Arizona, and lower default rates in states such as Idaho, Wyoming, North Dakota, and Nebraska. In Figure 10 in Appendix A.2, we additionally report the spatial default rates separately for every year. Spatial default patterns change visibly over time.
@@ 3.2 Predictor variables
From the data provided by Freddie Mac, we construct several static and temporally varying predictor variables. The latter predictor variables are based on information available at the beginning of each year and include the age of the loan in months, the current loan-to-value ratio, and the interest rate spread. The current loan-to-value ratio is calculated by dividing the current unpaid principal balance by the appraised value of the property at the time of origination. To calculate the interest rate spread, we follow Calabrese et al. (2024) and subtract from the loan interest rates the average interest rate currently issued by Freddie Mac for 30-year fixed-rate mortgages. This average interest rate is retrieved from the Federal Reserve Bank of St. Louishttps://fred.stlouisfed.org/series/MORTGAGE30US. Static predictor variables include the borrower’s credit score at the time when the mortgage originated, additional financial indicators, and several characteristics about the property and the mortgage. These predictor variables are commonly chosen when Freddie Mac’s loan-level data set is used in the credit risk literature; see, e.g., Hu and Zhou (2019) and Medina-Olivares et al. (2023a). In addition, we include a number of macroeconomic predictor variables at state and national levels, which vary over time and are primarily sourced from the U.S. Bureau of Economic Analysishttps://bea.gov. A description of all predictor variables is provided in Table 1. Missing values for categorical predictor variables are imputed using the most frequent category, and numeric predictor variables are imputed using the mean. Summary statistics of the numeric and categorical variables are provided in Tables 4 and 5, respectively, in Appendix A.3. For all linear models, we additionally include an intercept term in the predictor variables $X_{ki}$.
@@
orig_dept_to_income Original dept-to-income: borrower’s monthly debt payments divided by the borrower’s monthly income
orig_loan_to_value Original combined loan-to-value: original mortgage loan plus any secondary mortgage loan divided by the mortgage appraised value
orig_unpaid_balance Original unpaid principal balance of the mortgage
@@
curr_loan_to_value Current loan-to-value: current unpaid principal balance divided by the mortgage appraised value
@@
gdp_growth Annual gross domestic product (GDP) growth rate by state
income_per_capita Logarithmic per capita personal income by state
expend_per_capita Logarithmic per capita personal consumption expenditures by state
unemployment_rate Annual unemployment rate by state
hpi_growth Annual house price index growth rate by state
construction_growth Annual growth rate of GDP attributable to the construction industry by state
op_surplus_growth Annual growth rate of gross operating surplus across all industries by state
inflation_rate Annual U.S. inflation rate
djia_growth Annual growth rate of the year-end Dow Jones Industrial Average (DJIA)
@@ 3.3 Models considered and implementation details
We consider the following models with increasing levels of complexity: an independent linear hazard model, a linear spatial model, a linear spatio-temporal model, a tree-boosted spatial frailty model, and a tree-boosted spatio-temporal frailty model; see Section 2. We have also tried using an independent linear hazard model with dummy-coded fixed effects for the three-digit postal codes. However, the estimation of this model crashed due to memory overload with the R package stats and numerical instabilities with the R package biglm. For all models that include a spatial or spatio-temporal Gaussian process, we use a Matérn covariance function as defined in Equations (4) and (5) with smoothness parameter $\nu=1.5$ and $m=20$ nearest neighbors for estimation and prediction with Vecchia approximations. We do not estimate the smoothness parameter $\nu$ since “binary data contain no information about [$\nu$]” (De Oliveira, 2000). For purely spatial processes, we use the Euclidean distance to determine the nearest neighbors in Vecchia approximations and a random ordering of the spatial coordinates since this gives accurate approximations (Guinness, 2018). For the spatio-temporal models, we use a correlation-based approach to determine nearest neighbors as in Kang and Katzfuss (2023) and an increasing temporal order and a random spatial order for coordinates with the same time. Following Kang and Katzfuss (2023), the nearest neighbors for the spatio-temporal models are redetermined in every iteration that is a power of two whenever an optimization algorithm is used for determining the parameters $\theta$. For prediction with Vecchia approximations, the observed points appear first in the ordering, and the Gaussian process at prediction points is only conditioned on the training data in the Vecchia approximation. For finding optima for the parameters $\theta$, we use the limited-memory BFGS algorithm. Estimation and prediction with the linear spatial, linear spatio-temporal, tree-boosted spatial, and tree-boosted spatio-temporal frailty models is done using the GPBoost§§§https://github.com/fabsig/GPBoost library version 1.5.8 (Sigrist et al., 2021). The code for reproducing the mortgage credit risk application and generating the data set is publicly available at https://github.com/pkuendig/SpaceTimeML.
@@ 3.4 Sample split for model evaluation and choosing tuning parameters
All the following results are based on temporal out-of-sample predictions. Specifically, we conduct one-year-ahead default predictions for each year starting from 2008 through 2022 using an expanding window training data approach. In other words, we first consider all loan-year observations up to and including the year 2007 as training data and make predictions for all loan-year observations of 2008. We then continue by expanding the training window by one year and using the subsequent year as test data. Under this train-test split, the test data sets contain both new loans, which originated in the most recent year before the test data time period and are therefore not included in the training data, and older but still active loans, for which the training data contains historical observations.
For choosing tuning parameters, we analogously split every training window data set into an inner training data set and a validation data set. The validation data contains all loan-year observations of the most recent year in the training data, and the inner training data consists of all samples excluding this most recent year. Tuning parameters are chosen by estimating models on the inner training data and selecting the combinations of tuning parameters that maximize the area under the receiver operating characteristic curve (AUC) on the validation data. The candidate tuning parameters for the tree-boosted spatial and spatio-temporal frailty models are shown in Table 6 in Appendix A.4.
@@ 3.5 Prediction of individual default probabilities
Table 2 reports the AUC, H-measure, mean log-loss, Brier score, and ECE for every model, averaged over the 15 years for which we perform one-year-ahead default predictions. We find that the predictive default probabilities of the tree-boosted spatio-temporal frailty model are the most accurate in terms of AUC and H-measure, while predictive default probabilities of the purely spatial LaGaBoost model are the most accurate in terms of mean log-loss, Brier score, and ECE. Adding a spatial frailty variable to an independent linear hazard model improves the prediction accuracy. Compared to a purely spatial linear model, the linear spatio-temporal model has a higher prediction accuracy in terms of AUC and H-measure, but a lower accuracy in terms of mean log-loss, Brier score, and ECE. Overall, predictive default probabilities obtained using the tree-boosted spatial model are consistently more accurate than those obtained using the linear spatial model. The same holds true when comparing the tree-boosted spatio-temporal model and the linear spatio-temporal model. This suggests the presence of non-linearities and/or interactions among the predictor variables, a conclusion we corroborate through a more detailed analysis of the estimated models in Section 3.7.
@@
Linear independent 0.7676 0.2024 $7.926\times 10^{-2}$ $1.663\times 10^{-2}$ $1.128\times 10^{-2}$
Linear spatial 0.7702 0.2050 $7.924\times 10^{-2}$ $1.661\times 10^{-2}$ $1.112\times 10^{-2}$
Linear spatio-temporal 0.7711 0.2081 $8.124\times 10^{-2}$ $1.697\times 10^{-2}$ $1.342\times 10^{-2}$
LaGaBoost spatial 0.7716 0.2065 $\mathbf{7.816\times 10^{-2}}$ $\mathbf{1.633\times 10^{-2}}$ $\mathbf{8.993\times 10^{-3}}$
LaGaBoost spatio-temporal 0.7772 0.2149 $7.944\times 10^{-2}$ $1.658\times 10^{-2}$ $1.066\times 10^{-2}$
@@
Figure 3 additionally reports the AUC of every model over time. We observe that the AUCs are relatively high for all models and for all years before the COVID-19 pandemic. In particular, the AUC remains at high levels during the global financial crisis around the year 2008. However, in the year 2020 during the COVID-19 pandemic, the AUC drops to considerably lower levels for all models and increases again in the subsequent years. Similar patterns over time are observed for the H-measure and the ECE; see Figures 12 and 13 in Appendix A.6. We interpret this sudden decrease in prediction accuracy of all models in 2020 in the sense that an artificial, external shock in the form of lock-downs led to very different default mechanisms compared to normal times and also compared to the global financial crisis around the year 2008.
It is important to note that although the observed differences in AUC between the models are relatively small, even small improvements in prediction accuracy yield large economic gains, as the value of the mortgage portfolio of all active loans is in the billions of US dollars. Furthermore, we report in Table 8 in Appendix A.6 p-values of pairwise DeLong tests (DeLong et al., 1988) to assess whether the differences in AUCs of the pooled predictive default probabilities over all test years are significant across the different models. We find that all models have significantly different AUCs at the 5% level, except when comparing the linear independent and linear spatial models.
Figure 3: Temporal out-of-sample test area under the receiver operating characteristic curve (AUC) (higher = better).
In Table 7 in Appendix A.5, we also report the average wall clock times for training the different models on data up to the year 2014. The spatio-temporal models require more time compared to the spatial models due to the higher dimensionality of the space-time Gaussian processes. Furthermore, the tree-boosted frailty models are slower compared to their corresponding linear counterparts. In detail, the estimation times for the spatial and spatio-temporal tree-boosting models are approximately one and 30 minutes, respectively. Calculations were performed on a laptop with an Intel i7-12800H processor and 32 GB of random-access memory.
@@ 3.6 Prediction of loan portfolio loss distributions
Table 3 reports the CRPS, the 99% quantile loss, and the RMSE for the 15 years used as test data. Overall, the tree-boosted spatial and spatio-temporal frailty models are the most accurate models. In more detail, we observe that the tree-boosted spatio-temporal model has the lowest quantile loss. In other words, this model yields the most accurate predictive upper quantiles of portfolio loss distributions. Furthermore, the tree-boosted spatial model has the lowest CRPS and RMSE, i.e., this model is the most accurate for the center of portfolio loss distributions. In detail, the difference in RMSE between the tree-boosted spatial model and the linear spatial model are approximately US \$35 million. In addition, the predictive portfolio loss distributions of the linear spatial model are more accurate in terms of the CRPS and the RMSE compared to the independent linear hazard model.
@@
Linear independent $2.699\times 10^{8}$ $1.158\times 10^{8}$ $4.269\times 10^{8}$
Linear spatial $2.627\times 10^{8}$ $1.210\times 10^{8}$ $4.178\times 10^{8}$
Linear spatio-temporal $3.128\times 10^{8}$ $1.188\times 10^{8}$ $5.615\times 10^{8}$
LaGaBoost spatial $\mathbf{2.034\times 10^{8}}$ $1.072\times 10^{8}$ $\mathbf{3.823\times 10^{8}}$
LaGaBoost spatio-temporal $2.598\times 10^{8}$ $\mathbf{1.021\times 10^{8}}$ $4.960\times 10^{8}$
@@
Next, Figure 4 shows the differences between the means of the predictive portfolio loss distributions and the realized portfolio losses over the years 2008 to 2014. We see that the tree-boosted spatial and spatio-temporal frailty models clearly predict the most realistic mean portfolio losses during the global financial crisis around the year 2008. The linear models clearly overestimate the portfolio loss in the second year of the global financial crisis, 2009, a trend that is reversed by underestimating it in 2010. In Figure 14 in Appendix A.7, we also show the differences between the predictive mean portfolio losses and the realized portfolio losses for all prediction years from 2008 to 2022.
@@
Figure 5 additionally shows the time series of predictive $99\%$ quantiles of one-year-ahead portfolio loss distributions. We again find that the tree-boosted frailty models predict the most consistent upper tail loan portfolio losses at the beginning and end of the global financial crisis. Similar to the mean portfolio loss, linear models initially overestimate, then underestimate, and finally overestimate upper tail losses during the years of the financial crisis. The linear independent model without a frailty Gaussian process generates the smallest predictive $99\%$ quantiles before the global financial crisis in 2008. The sudden one-year default spike in 2020 during the COVID-19 pandemic causes all models to predict very high upper tail loan portfolio losses for the subsequent year 2021. However, the overestimation of the upper tail portfolio loss in Figure 5 as well as the mean portfolio loss in Figure 14 in Appendix A.7 in 2021 is particularly pronounced for the linear spatio-temporal model and the tree-boosted spatio-temporal model. Purely spatial frailty models, on the other hand, yield lower predictive quantiles for the year 2021 since their frailty components are constant over time.
@@ 3.7 Model interpretation
In the following, we aim to better understand the functioning of the tree-boosted and linear spatio-temporal frailty models. In Figure 6, we show the mean of the posterior distribution of the latent spatio-temporal frailty Gaussian process for the tree-boosted spatio-temporal model when training on data up to and including the year 2021, which corresponds to the model for predicting defaults in the most recent year 2022. Due to space constraints, posterior means are not shown for the year 2000. We additionally show these posterior mean maps for the linear spatio-temporal model in Figure 15 in Appendix A.8. We observe that there is considerable variation in the latent frailty variable over space and this variation changes over time. In particular, for the years 2005 and 2006, posterior means for the latent Gaussian variable are high in the New Orleans area, which can be explained by high default rates following Hurricane Katrina in 2005. In 2008 and subsequent years, posterior means are high in the so-called “bubble” states California, Florida, Nevada, and Arizona, where house prices rose particularly rapidly in the run-up to the subprime mortgage crisis (Haughwout et al., 2011). For the year 2020, posterior means are high due to the generally large number of defaults, with the highest values found in the New York City area (not visible at this resolution). This is not surprising given that New York City was one of the first and most affected areas in the U.S. by the COVID-19 pandemic. Comparing posterior means of the tree-boosted spatio-temporal frailty model with those of the linear spatio-temporal model, we find similar patterns over space and time.
@@
In Figure 7, we report the estimated covariance parameters for the different expanding window training data sets for the linear spatial, the linear spatio-temporal, the tree-boosted spatial, and the tree-boosted spatio-temporal frailty models. For the latter, estimated variance and range parameters are relatively constant until the year 2020, when the variance parameter increases from 0.45 to 1.13, and in the following year 2021, the range parameter for time decreases from 2.58 to 1.08. These estimates are likely a consequence of the sudden atypical one-year default spike in 2020 during the COVID-19 pandemic. Since default mechanisms were likely very different in this year compared to the past due to an artificial external shock, less variation in the response variable is explained by the fixed effects and the latent Gaussian variable becomes more important as can be inferred from the higher marginal variance. Subsequently, defaults return to normal levels in 2021, and the estimated correlation with previous years is lower when including 2020 in the training data as can be seen from the lower estimated temporal range parameter. In both the linear spatial model and the tree-boosted spatial model, the estimated range parameters decrease during the initial training years and then remain relatively constant for the remaining years. Furthermore, the estimated variance parameter of the linear spatial model is consistently higher than the one of the tree-boosted spatial model throughout all training years. This indicates that the latent frailty variable plays a more important role in the linear spatial model, since its linear deterministic function can explain less variation in the data compared to the tree-ensemble of the LaGaBoost spatial model. Similarly, for training years 2011 to 2019, the estimated variance parameter of the linear spatio-temporal model is also consistently higher than that of the LaGaBoost spatio-temporal model. This difference can also be attributed to the different explanatory power of their respective predictor functions.
@@
For understanding the function $F(\cdot)$ of the tree-boosted frailty models, we use SHAP values and SHAP dependence plots (Lundberg and Lee, 2017). Specifically, we consider SHAP values for the tree-boosted spatio-temporal model trained on data up to the year 2014. We chose this training window because, for the corresponding prediction year 2015, there are pronounced differences in the accuracy of predictive default probabilities between the tree-boosted and linear frailty models. SHAP values are calculated using 10’000 randomly selected instances of the training data. This subsampling is done to reduce the computational complexity and to avoid that SHAP dependence plots are overcrowded. Using a larger random subsample or a different subsample gives almost identical results (results not shown). Figure 8 shows the SHAP values. The predictor variables are ordered according to the average of the absolute values of the SHAP values. According to these results, the eight most important predictor variables in descending order are the credit score (credit_score), the interest rate spread (ir_spread), the current loan-to-value (curr_loan_to_value), the indicator whether multiple borrowers are obligated to repay (multiple_borrowers), the age of the loan (n_months), the indicator whether the mortgage is used for purchase (loan_purpose = P), the original debt-to-income (orig_dept_to_income), and the original unpaid principal balance (orig_unpaid_balance). For the six numeric variables thereof, we report the corresponding SHAP dependence plots in Figure 9. In all dependence plots, we observe strong interaction effects. This can be inferred from the large vertical scatter of the SHAP values and the systematic colored relationship with other predictor variables. For example, the slope of the SHAP values for the credit score variable (credit_score) is less negative if the interest rate spread variable (ir_spread) is positive. I.e., when interest rate spreads are high, the individual credit score matters less compared to when interest rate spreads are low. Furthermore, we find strong non-linear effects for the age of the loan (n_months) and the original unpaid principal balance (orig_unpaid_balance). For instance, the age of a loan is positively related to the default probability up to a certain age of approximately three years, and then the effect flattens out. In addition, the variables interest rate spread (ir_spread), current loan-to-value (curr_loan_to_value), and original debt-to-income (orig_dept_to_income) also have clearly non-linear relationships with slopes that change markedly for large and small values of these variables. For instance, increasing interest rate spreads are related to higher default probabilities, but this effect only holds up to a certain level of approximately two and a half percent, after which the relationship levels off, and additional increases in interest rate spreads are not associated with higher default probabilities. Furthermore, the debt-to-income (orig_dept_to_income) appears to have an approximately logistic-shaped effect on the default probability with the relationship being almost flat for both small debt-to-income values below 20 percent and large values above 50 percent, and in between, the effect of debt-to-income is approximately linear with a slope that interacts, among other things, with the loan age. We conclude that there are strong interaction and non-linear effects. This is likely also the reason why the tree-boosted spatial and spatio-temporal frailty models achieve higher prediction accuracy than their respective linear counterparts.
@@
Figure 8: SHAP values for the tree-boosted spatio-temporal frailty model when training on data up to the year 2014.
@@
Figure 9: SHAP dependence plots for the tree-boosted spatio-temporal frailty model when training on data up to the year 2014.